help me find an equation of a line containing the given point and having the given slope. (4,9),m=15/8
Answer by lwsshak3(11628)

help me find an equation of a line containing the given point and having the given slope. (4,9),m=15/8
***
Standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
Equation of given line: y=15x/8+b
solve for b using coordinates of given point (4,9)
9=15*4/8+b
9=60/8+b
b=72/8-60/8=12/8=3/2
equation of given line: y=(15/8)x+3/2